OpenStudy (trisarahtops):

Use the table below to evaluate d/dx [g[f(2x]] at x=1. x 1 2 3 4 f(x) 6 1 8 2 f ′(x) 1 3 5 7 g(x) 1 4 4 3 g ′(x) 4 5 5 –4

OpenStudy (solomonzelman):

Use the table below to evaluate d/dx [g[f(2x)] at x=1. \(\bf ~~x \quad \quad~ 1\quad \quad~ 2\quad ~~~3\quad~~~4\) `f(x) 6 1 8 2` f ′(x) 1 3 5 7 `g(x) 1 4 4 3` g ′(x) 4 5 5 –4 making it more clear

OpenStudy (solomonzelman):

Apply the chain rule to differentiate the g.

OpenStudy (trisarahtops):

wait don't we have to set it up fiirst

OpenStudy (solomonzelman):

you already have the set up

OpenStudy (solomonzelman):

You need to evaluate g'(1) .... no?

OpenStudy (solomonzelman):

Isn't that the meaning of evaluating the derivative (d/dx) of g(x) at x=1??

OpenStudy (solomonzelman):

So do you agree that you need g'(1) ?

OpenStudy (trisarahtops):

uh I guess not sure. Idk i'm just confused on how to start

OpenStudy (solomonzelman):

The derivative of \(\color{#000000 }{ \displaystyle f(x) }\) is \(\color{#0000ff }{ \displaystyle f'(x) }\). However, if you have a function of \(x\) inside the \(f(x)\), then The derivative of \(\color{#000000 }{ \displaystyle f(g(x)) }\) is NOT just \(\color{#ff0000}{ \displaystyle f'(g(x)) }\). Rather, the derivative of \(\color{#000000 }{ \displaystyle f(g(x)) }\) is (YES) \(\color{#0000ff }{ \displaystyle f'(g(x))\cdot g'(x) }\), and we get this by applying the Chain Rule. And what happens when you have another function of x inside the g(x)- (in your case "2x"), you apply the chain rule once again. The derivative of \(\color{#000000 }{ \displaystyle f(g(2x)) }\) is NOT \(\color{#ff0000 }{ \displaystyle f'(g(2x))\cdot g'(2x) }\); Rather, the derivative of \(\color{#000000 }{ \displaystyle f(g(2x)) }\) is \(\color{#0000ff }{ \displaystyle f'(g(2x))\times g'(2x) \times 2 }\)

OpenStudy (solomonzelman):

I might have said in a too lengthy way, but I just want to make sure you understand that the chain rule is applied multiple times.

OpenStudy (trisarahtops):

no no this makes more sense to me. so since x=1 it's just f'(g(2)) x g'(2) x 2

OpenStudy (solomonzelman):

yes, very good

OpenStudy (trisarahtops):

so now i plug in what's in the chart

OpenStudy (solomonzelman):

yes

OpenStudy (trisarahtops):

f'(4)) x 5 x 2 7 x 10 70

OpenStudy (solomonzelman):

Yes, fabulous !!

OpenStudy (trisarahtops):

Thank you so much ;)

OpenStudy (solomonzelman):

Anytime!

OpenStudy (ayeedomo):

It's domo! :D

OpenStudy (solomonzelman):

domo?

OpenStudy (ayeedomo):

domo icon

OpenStudy (trisarahtops):

yup. someone in a domo suit ;D

OpenStudy (ayeedomo):

lol

OpenStudy (solomonzelman):

.... \( ; ) \)